OpenStudy (anonymous):

If a mass of 2.0 kg is lifted through a height of 5.0 m, what force does it require to lift it at constant speed?

OpenStudy (anonymous):

mg

OpenStudy (anonymous):

Doesn't matter how high its lifted. The gravitational force experience = mass * acc. due to gravity. To balance this, an equal force in the opposite direction is required. So first calculate the gravitational force

OpenStudy (anonymous):

@amishra

OpenStudy (anonymous):

Gravitational force is 9.81 m/s^2 ? If so, Then we just use f=ma like F= 2*9.81 ? Is this correct?

OpenStudy (anonymous):

Wait... your question is what force is needed to keep it stationary or at constant speed. If stationary...do what I told you If constant speed, you need a net force of 0

OpenStudy (anonymous):

but yeah that's correct

OpenStudy (anonymous):

Wait, what force is needed to lift the object, assuming that we are moving at constant speed. So the force should come out to be negative right?

OpenStudy (anonymous):

Lift the object at constant speed..? Force is a vector unit so depending on which direction you put it could be pos or neg

OpenStudy (anonymous):

That's the question :/ Okay, thank you so much! :D

OpenStudy (anonymous):

sure. anytime